# OpenECAD: An Efficient Visual Language Model for Computer-Aided Design

### v1

@@ Abstract
Computer-aided design (CAD) tools are utilized in the manufacturing industry for modeling everything from cups to spacecraft. These programs are complex to use and typically require years of training and experience to master. Structured and well-constrained 2D sketches and 3D constructions are crucial components of CAD modeling. A well-executed CAD model can be seamlessly integrated into the manufacturing process, thereby enhancing production efficiency. Deep generative models of 3D shapes and 3D object reconstruction models has garnered significant research interest. However, most of these models are represented in discrete forms. Moreover, the few models based on CAD operations often have substantial input restrictions. In this work, we fine-tuned pre-trained models to create OpenECAD (0.55B, 0.89B, and 4.2B), leveraging the visual, logical, coding, and general capabilities of visual language models. OpenECAD can process images of 3D designs as input and generate highly structured 2D sketches and 3D construction commands. These outputs can be directly used with existing CAD tools’ APIs to generate project files. To train our network, we created a new CAD dataset. This dataset is based on existing public CAD datasets, with adjustments and augmentations to meet the requirements of  VLM training.
Index Terms:
@@
I Introduction
@@
A typical 3D part drawing process involves multiple ”Sketch-Extrusion” steps. The relationship between sketch and extrusion is not strictly one-to-one; a single sketch can correspond to multiple extrusions, and one extrusion can be composed of multiple sketches. A sketch consists of multiple closed loops formed by several lines and includes constraints within or between them to ensure the sketch is fully defined. The extrusion operation generates a 3D feature based on the 2D sketch(s). Generally, there is a sequential relationship between ”Sketch-Extrusion” steps, as a sketch’s reference plane may rely on an existing face in the model, and constraints within the Sketch may depend on points or lines already present in the model. As shown in Figure 1, this is a simple process for drawing 3D parts.
@@
To address this issue, some studies have emerged with the development of neural networks. Since 3D shapes are still based on 2D sketches, these studies can be divided into 2D sketch generation and 3D CAD model generation. SketchGraphs[18] provides a dataset of constrained 2D sketches, CurveGen and TurtleGen[19] generate closed curve sketches, and CADL[20] uses language models to generate sketches. For 3D model generation, SolidGen[21] and ComplexGen[22] generate B-rep models, while Point2Cyl[23], DeepCAD[24], and Free2CAD[25] generate CAD command sequences. It is worth noting that these models often have significant input limitations. For example, they may require existing 3D point clouds, may not accept inputs to limit how to generate, or require isometric hand-drawn sketches. These constraints still differ significantly from existing manual modeling workflows, making it challenging to apply them in production processes.
@@
    Based on existing public CAD datasets, such as SketchGraphs[18], Fusion 360 Gallery[26], and DeepCAD[24], we convert these datasets into ”image-CAD operation code” pairs, including 2D sketch generation and 3D shape generation. Using currently available multimodal large language models, such as GPT-4o, we generate high-quality annotated datasets. Additionally, we utilize rendering tools to generate various forms of images to enhance the versatility of model inputs.
@@
    Based on existing public small language models and multimodal small language models, such as TinyLLaMA[27], OpenELM[28], Phi3[29], and TinyLLaVA[30], we train them using the aforementioned training sets and test their performance in CAD design.
@@
    Design a program standard for CAD operations that can interface with the APIs of various CAD tools, enabling the generated model code to be directly used in CAD applications.
II Related Work
II-A Small Language and Visual Language Models
Small language models have recently become very popular due to the growing demand for privacy and local execution. Small language models are those that are smaller in size and require fewer computational resources, allowing them to run on local devices and ensure data privacy and security. Despite their smaller size, these models can still deliver efficient performance on specific tasks, making them suitable for resource-constrained environments. Currently available small language models include Microsoft’s Phi series[29], Apple’s OpenELM[28] and TinyLLaMA[27], based on Meta’s LLaMA[31].
@@ II-A Small Language and Visual Language Models
II-B Generative models of 3D shapes’ CAD commands
@@ III Method
III-A Design and Generation of Datasets
@@
III-A1 Definition of the Code Format for CAD Operation Sequences
Mature CAD tools support a rich set of commands, most of which are simplified calls to collections of basic commands—for example, a rectangle is a collection of four line segments and some constraints. However, in practice, only a small subset of these commands is widely used. Therefore, we currently consider only the most basic and commonly used commands in our dataset (see Table I). Due to the versatility of language models, extending the dataset later is relatively easy.
TABLE I: CAD commands and their parameters.

### v2

@@ Abstract
Computer-aided design (CAD) tools are utilized in the manufacturing industry for modeling everything from cups to spacecraft. These programs are complex to use and typically require years of training and experience to master. Structured and well-constrained 2D sketches and 3D constructions are crucial components of CAD modeling. A well-executed CAD model can be seamlessly integrated into the manufacturing process, thereby enhancing production efficiency. Deep generative models of 3D shapes and 3D object reconstruction models have garnered significant research interest. However, most of these models are represented in discrete forms. Moreover, the few models based on CAD operations often have substantial input restrictions. In this work, we fine-tuned pre-trained models to create OpenECAD (0.55B, 0.89B, 2.4B, 3.2B and 4.2B), leveraging the visual, logical, coding, and general capabilities of visual language models. OpenECAD can process images of 3D designs as input and generate highly structured 2D sketches and 3D construction commands. These outputs can be directly used with existing CAD tools’ APIs to generate project files. To train our network, we created a new CAD dataset. This dataset is based on existing public CAD datasets, with adjustments and augmentations to meet the requirements of VLM training.
keywords:
@@
journal: Computer-Aided Design
[…]
1 Introduction
@@
A typical 3D part drawing process involves multiple ”Sketch-Extrusion” steps. The relationship between sketch and extrusion is not strictly one-to-one; a single sketch can correspond to multiple extrusions, and one extrusion can be composed of multiple sketches. A sketch consists of multiple closed loops formed by several lines and includes constraints within or between them to ensure the sketch is fully defined. The extrusion operation generates a 3D feature based on the 2D sketch(es). Generally, there is a sequential relationship between ”Sketch-Extrusion” steps, as a sketch’s reference plane may rely on an existing face in the model, and constraints within the Sketch may depend on points or lines already present in the model. As shown in Figure 1, this is a simple process for drawing 3D parts.
@@
To address this issue, some studies have emerged with the development of neural networks. Since 3D shapes are still based on 2D sketches, these studies can be divided into 2D sketch generation and 3D CAD model generation. SketchGraphs[18] provides a dataset of constrained 2D sketches, CurveGen and TurtleGen[19] generate closed curve sketches, and CADL[20] uses language models to generate sketches. For 3D model generation, SolidGen[21] and ComplexGen[22] generate B-rep models, while Point2Cyl[23], DeepCAD[24], and Free2CAD[25] generate CAD command sequences. It is worth noting that these models often have significant input limitations. For example, they may require existing 3D point clouds, may not accept inputs to limit how to generate, or require isometric hand-drawn sketches. These constraints significantly differ from existing manual modeling workflows, making it challenging to apply them in production processes.
@@
    We established OpenECAD dataset. Based on existing public CAD datasets, such as SketchGraphs[18], Fusion 360 Gallery[26], and DeepCAD[24], we convert these datasets into ”image-CAD operation code” pairs, including 2D sketch generation and 3D shape generation. Using currently available multimodal large language models, such as GPT-4o, we generate high-quality annotated datasets. Additionally, we utilize rendering tools to generate various forms of images to enhance the versatility of model inputs.
@@
    We trained OpenECAD model. Based on existing public small language models and multimodal small language models, such as TinyLLaMA[27], OpenELM[28], Gemma[29], Phi-2, Phi-3-mini[30], and TinyLLaVA[31], we train them using the aforementioned training sets and test their performance in CAD design.
@@
    We integrated the code generated by OpenECAD into CAD tools. We design a program that can interface with the APIs of various CAD tools, enabling the generated model code to be directly used in CAD applications.
2 Related Work
2.1 Small Language and Visual Language Models
Small language models have recently become very popular due to the growing demand for privacy and local execution. Small language models are those that are smaller in size and require fewer computational resources, allowing them to run on local devices and ensure data privacy and security. Despite their smaller size, these models can still deliver efficient performance on specific tasks, making them suitable for resource-constrained environments. Currently available small language models include Microsoft’s Phi series[30], Apple’s OpenELM[28], Google’s Gemma[29] and TinyLLaMA[27], based on Meta’s LLaMA[32].
@@ 2.1 Small Language and Visual Language Models
2.2 Generative models of 3D shapes’ CAD commands
@@ 3 Method overview
4 Generation of Datasets
In this chapter, we encompass several key components essential for OpenECAD datasets. We cover the design and generation of datasets (Section 4.1), the definition of the code format for CAD operation sequences (Section 4.2), and the translation of these sequences (Section 4.3). Additionally, we address the generation of natural language descriptions for CAD models (Section 4.4), the production of CAD model view images (Section 4.5), and the overall creation of the CAD dataset (Section 4.6).
4.1 Design and Generation of Datasets
@@
4.2 Definition of the Code Format for CAD Operation Sequences
Mature CAD tools support a rich set of commands, most of which are simplified calls to collections of basic commands—for example, a rectangle is a collection of four line segments and some constraints. However, in practice, only a small subset of these commands is widely used. Therefore, we currently consider only the most basic and commonly used commands in our dataset (see Table 1). Due to the versatility of language models, extending the dataset later is relatively easy.
@@ 4.2 Definition of the Code Format for CAD Operation Sequences
Sketch and its helper
add_sketchplane origin_point, x_axis, y_axis
add_profile loops_list
add_sketch sketchplane, profile
Extrusion
add_extrude sketch, operation, type, extent_size
Table 1: CAD commands and their parameters.
Curves. In practice, we consider three types of widely used curve commands: drawing a line, drawing an arc, and drawing a circle. The line is defined by its starting and ending points, the arc is determined by its starting point, ending point, and midpoint, and the circle is specified by its center and radius.
All operation names are chosen to be as meaningful as possible to aid the language model in understanding and using them. With these operation commands, we describe a CAD model as a sequence of executable commands in code. The pseudo code corresponding to a pair of Sketch-Extrusion operations is shown in the Algorithm 1.
Algorithm 1 Pseudo Code Example
$SketchPlane$ = add_sketchplane($origin$, $normal$, $x_{axis}$, $y_{axis}$)
$Loops$ and $Curves$ are lists.
$Line1$ = add_line($start$, $end$)
$Line1$ is appended to $Curves$.
$Loop$ = add_loop($Curves$)
$Loop$ is appended to $Loops$.
$Profile$ = add_profile($Loops$)
$Sketch$ = add_sketch($SketchPlane$, $profile$, $position$, $size$)
$Extrude$ = add_extrude($Sketch$, $operation$, $type$, $extent$)
4.3 Translation of CAD Operation Sequences
According to the definitions provided in the previous section, we can convert any CAD model covered by the described operations (in Table 1) into a CAD operation sequence as defined by us. However, in practice, different CAD tools or datasets may have slight variations in their definitions of CAD operations (see Figure 4). Therefore, we determine and transform each curve-drawing operation according to its degrees of freedom to match the syntax of the OpenECAD dataset. We have developed scripts to convert other CAD models or datasets into CAD operation sequence codes for training our VLM.
4.4 Generation of Natural Language Descriptions for CAD Models
Describing CAD models in natural language is crucial for understanding 3D shapes. Adding these descriptions as annotations to the CAD operation sequence code can help people quickly grasp the meaning of each code segment and swiftly locate and modify the necessary parts. To construct natural language descriptions corresponding to CAD operation sequence codes, we employed existing large language models such as GPT-4o. By using prompts and examples, we added natural language annotations to the existing CAD operation sequence codes. The prompt used was:
“Please add comments to the drawing code below to indicate what shape is drawn by each section of the code. Don’t add comment for every line or arc.”
To prevent the language model from adding annotations to every CAD operation (like drawing lines or arcs), we constrained the model to add annotations by operation units, such as drawing a rectangle, etc.
4.5 Generation of CAD Model View Images
4.6 Creation of CAD Dataset
Based on the above method, we generated a 3D CAD dataset derived from DeepCAD dataset[24] and Fusion 360 Gallery Reconstruction Dataset[26]. We excluded designs that were generated incorrectly or had overly long CAD command sequences due to the number of tokens that small language models can accept is limited, resulting in a total of 130,239 3D designs. The statistical distribution of the number of ”Sketch-Extrusion” steps in the 3D dataset is shown in the Figure 5.
5 Training and Fine-tuning Vision Language Models
In this chapter, we introduce the process of obtaining OpenECAD models. We discuss the choice of language models and multimodal approach (Section 5.1), as well as the fine-tuning of a pre-trained visual language model (Section 5.2).
5.1 The choice of language models and multimodal approach
Due to the high confidentiality requirements of manufacturing enterprises and the substantial CPU and GPU resources typically occupied by CAD design tools, we chose OpenELM[28] (450M), Gemma[29] (2B), Phi-2 (2.7B) and Phi-3 mini[30] (3.8B) with relatively small parameter sizes as the foundation for our language model. OpenELM is an efficient language model family developed by Apple. It employs a layer-wise scaling strategy to allocate parameters efficiently within each layer of the transformer model. This approach enhances accuracy while maintaining computational efficiency. Gemma is a family of lightweight, text-to-text, decoder-only large language models developed by Google. Phi, developed by Microsoft Research, is another series of advanced small language models (SLMs). This series includes Phi-2 (2.7B), Phi-3-mini (3.8B), Phi-3-small (7B), and Phi-3-medium (14B), each with different parameter sizes. Despite their compact parameter scales, Phi-3 models achieve language understanding and reasoning capabilities comparable to larger models. These models are trained on high-quality data and optimized algorithms. It is worth noting that the context acceptance length for OpenELM (450M) is 2048 tokens. In contrast, Gemmma and Phi has a minimum context acceptance length of 3072 tokens, with an option to use a version supporting up to 128k tokens at the cost of some content accuracy.
To endow the language model foundation with multimodal capabilities, we selected OpenAI’s CLIP[34] and Google’s SigLIP [35] as visual encoders for OpenELM (450M), Gemma (2B) and Phi-2 (2.7B). These models are trained on large datasets consisting of images paired with corresponding text descriptions, combining visual and textual information to make predictions.
5.2 Fine-tuning on a pre-trained visual language model
OpenECAD 0.55B & 0.89B
CPU Intel Core i9-7960X
GPU Nvidia GeForce GTX 1080 Ti
RAM DDR4-2666 32 GB
OS Ubuntu Server 22.04 LTS
OpenECAD 2.4B & 3.1B
CPU AMD EPYC 9754 Processor (18 vCPU)
GPU 2 x Nvidia GeForce RTX 4090D
RAM 60 GB
OS Ubuntu Server 22.04 LTS
OpenECAD 4.2B
CPU AMD EPYC 9754 Processor (18 vCPU)
GPU Nvidia GeForce RTX 4090D
RAM 60 GB
OS Ubuntu Server 22.04 LTS
Table 2: Development and running environment.
To enable the general visual-language model to generate CAD code, we fine-tuned it using the LoRA[37] method on the pre-trained OpenELM-CLIP, OpenELM-SigLIP, Gemma-SigLIP and Phi-2-SigLIP models. LoRA stands for Low-Rank Adaptation of Large Language Models. In the field of natural language processing (NLP), there is a common paradigm that involves large-scale pre-training on general domain data followed by adaptation to specific tasks or domains. However, as we pre-train larger models, fully fine-tuning all model parameters becomes less feasible due to computational costs. LoRA proposes a novel approach. Instead of full fine-tuning, LoRA freezes the pre-trained model weights and injects trainable rank decomposition matrices into each layer of the Transformer architecture. This significantly reduces the number of trainable parameters for downstream tasks.
We performed LoRA fine-tuning on the pre-trained OpenELM-CLIP (0.55B) and OpenELM-SigLIP (0.89B) models using an Nvidia GeForce GTX 1080 Ti (11 GB VRAM). The Rank and Alpha for LoRA were set to 128 and 256, respectively. The learning rate is $10^{-4}$. Training for one epoch with our 3D dataset took approximately 60 hours. Through fine-tuning, we obtained the OpenECAD 0.55B and 0.89B models.
Similarly, we used 2 Nvidia GeForce RTX 4090 D (24 GB VRAM) to perform LoRA fine-tuning on the Gemma-SigLIP (2.4B) and Phi-2-SigLIP (3.1B) model. The Rank and Alpha for LoRA were set to 128 and 256, respectively. The learning rate is $10^{-4}$. Training for one epoch with our 3D dataset took approximately 14 hours. Through fine-tuning, we obtained the OpenECAD 2.4B and 3.1B models.
We used an Nvidia GeForce RTX 4090 D (24 GB VRAM) to perform LoRA fine-tuning on the Phi-3 Vision (4.2B) model. The LoRA rank and alpha were set to 256 and 512, respectively. The learning rate is $10^{-4}$. Training with our dataset for one epoch took approximately 20 hours. Through fine-tuning, we obtained OpenECAD 4.2B model.
The specific development, training, and runtime environments are shown in the Table 2. The loss curves for the OpenECAD 0.55B, 0.89B, 2.4B, 3.2B and 4.2B models during training are illustrated in the Figure 6. Due to the different batch sizes used during training, the loss curves are proportionally scaled. As shown, the models converge after a single round of fine-tuning.
Figure 6: The loss curves for the OpenECAD 0.55B, 0.89B, 2.4B, 3.2B and 4.2B models during training.
6 Experiments
In this chapter, we evaluate the ability of OpenECAD models to generate CAD designs based on input images and analyze the results. As there are no existing CAD models that use images for CAD generation, we first define evaluation metrics (Section 6.1) and create CAD designs for testing (Section 6.2). Following this, we test the models, obtain the results (Section 6.3), and analyze them (Section 6.4).
6.1 Evaluation Metrics
Figure 7: Example illustrations of CAD designs for evaluating.
Scoring Item Score (out of 100)
Is the code executable? 10
Accuracy of 3D shape construction 90
For each Loop-Extrusion pair (out of 100)
- 2D sketch positioning 20
- Accuracy of 2D Loop 60
- 3D operation 20
Table 3: Scoring Table.
6.1.1 Scoring Algorithm for Evaluating Outputs
We score the results based on executability and the correctness of the 3D shapes (see Table 3). Each Loop-Extrusion pair is considered a basic 3D shape. If the number of generated Loop-Extrusion pairs is insufficient, points will be deducted proportionally. For each Loop-Extrusion pair, we score based on the reference plane position of the 2D sketch, the extrusion operation, and the drawing of the 2D loop. The specific calculation method is shown in the following formula.
$score=10R+\frac{90}{N_{s}}\sum_{i=1}^{N_{s}}(20P_{i}+20O_{i}+60L_{j})$ (1)
6.2 CAD Designs for Test
6.3 Evaluation Results
6.3.1 The Inference Speed of the Models
For OpenECAD 0.55B and 0.89B, we used an Nvidia GeForce GTX 1080 Ti (11 GB VRAM) for inference, with a peak memory usage of less than 5GB. Getting a 3D shape’s code took approximately 20 seconds. For OpenECAD 2.4B, we used an Nvidia GeForce RTX 4090 D (24 GB VRAM) for inference, with a peak memory usage of less than 9GB. For OpenECAD 3.2B and 4.2B, we used an Nvidia GeForce RTX 4090 D (24 GB VRAM) for inference, with a peak memory usage of less than 18GB. Getting a 3D shape’s code took approximately 45 seconds.
6.3.2 Generating Code and Rendering
We tested OpenECAD versions 0.55B, 0.89B, 2.4B, 3.2B and 4.2B, and also tested GPT-4o. Some successfully rendered test results are illustrated in Figure 8. Since GPT-4o was not trained specifically for the use case described in this paper, we used the following prompt for testing:
”I am going to provide you with a 3D shape’s corresponding Python code for its modeling process. After that, I will input images of other 3D shapes, and please provide the corresponding Python code for them. $<$OpenECAD style Code Example$>$ This image is a view of a 3D model from a certain angle. Please write the codes of it in the format above.”
During the testing of OpenECAD, the maximum number of new tokens that 0.55B, 0.89B, 3.2B and 4.2B models can generate is 1536, 1024, 2048, 3072 and 3072, respectively. Due to these context limitations, there are instances where the generated code is incomplete. GPT-4o, due to its support for a longer context, did not encounter issues with incomplete code generation.
Similar to generating datasets, we use the PythonOCC tool to write a library that attempts to directly run the generated code to create CAD projects and output them as STEP files (B-Rep format). Then, we use PythonOCC to render the STEP files to obtain views. Since the generated programs from the model cannot ensure complete accuracy, errors may occur that the API operator cannot handle, leading to model rendering failures.
Based on the scoring method mentioned above, we obtained the scores, the number of executable results, the number of incomplete results, the number of runtime errors, and the number of completely correct results for each model, as shown in the Table 4.
Model Overall Executable Unfinished Run with Completely
Score Errors Correct
0.55B 46.52 36 21 0 15
0.89B 37.19 22 35 0 19
2.4B 59.86 50 7 0 20
3.1B 55.39 53 0 4 16
4.2B 51.91 51 4 2 14
GPT-4o 15.44 52 0 5 4
Table 4: Evaluation Results Table.
As shown in Table 4, OpenECAD 2.4B achieved the highest overall score and the most fully correct outputs. Both 0.55B and 0.89B models produced a significant amount of incomplete code. The 0.89B model’s higher token consumption for image inputs reduced the number of new tokens it could generate, leading to more incomplete outputs compared to the 0.55B model, and thus a lower overall score. The 3.1B and 4.2B models achieved relatively high scores, but the number of fully correct generations was lower than that of the 0.55B and 0.89B models. All OpenECAD models demonstrated significant advantages over the unrefined GPT-4o, whose sole advantage was its ability to handle the entire code due to its extended context length.
6.4 Analysis of OpenECAD’s outputs
6.4.1 Code Generation Analysis
For the generated code, the language model effectively adheres to the OpenECAD dataset’s syntax, even without being explicitly instructed on the specific details of the syntax. Regarding the curves in the sketch, the language model successfully ensures that consecutive curves are connected end-to-end, with the first and last curves also joined, thus maintaining closure. Additionally, the language model adeptly manages the positions of points to ensure the profile’s validity. For the fully generated code, the vast majority can be executed and rendered directly. For generated code with runtime errors, this includes both execution crashes and issues where the B-Rep file fails to render correctly.
6.4.2 Generation Error Analysis
Figure 9: Typical errors in generated results.
  • 1.
    Case 1 and Case 2: For complex and irregular shapes, the corresponding sketches often fail to generate correctly, or the profiles of the 2D sketches are inaccurately created. Possible reasons include the vision tower’s low resolution, which hinders accurate recognition of complex shapes, and the language model’s inability to comprehend complex shapes from the CAD operation code. These issues ultimately result in incorrect loops or intersecting loops, leading to erroneous sketch profiles and broken extrusions.
  • 2.
    Case 3 and Case 7: The model lacks spatial reasoning capabilities, resulting in incorrect handling of the size proportions between sketches and extrusions. This leads to extrusions that are either too long or too small, causing discrepancies in size relative to other parts.
  • 3.
    Case 4: When multiple ”Sketch-Extrusion” steps are involved, the positions of the sketches in subsequent steps are often incorrect, leading to errors in the final overall model.
  • 4.
    Case 5: Due to the minimal color difference between the outer surface of a solid cylinder and the inner wall of a hollow cylinder, the primary distinguishing feature is the inner wall circle at the top. This makes it difficult for the model to differentiate between solid and hollow cylinders, leading to confusion. This issue also extends to other similar shapes with holes.
  • 5.
    Case 6: The difficulty in distinguishing concave and convex shapes results in erroneous models. This issue may arise because the vision tower can detect these shapes but cannot differentiate between them, or because the language model fails to generate the correct code, leading to sketch errors.
  • 6.
    Case 8: Missing part(s). For some relatively subtle features, the model struggles to accurately identify them, resulting in missing features.
  • 1.
    The model has weak handling capabilities for complex sketches. This includes issues such as the vision tower’s low resolution, which leads to an inability to recognize shapes, and the language model’s failure to correctly process complex shapes, resulting in erroneous sketch generation.
  • 2.
    The generated models have poor size accuracy, particularly in handling the proportions between sketch dimensions and extrusion sizes. This results in mismatched parts, as the model fails to ensure consistent sizes between the sketches and extrusions of different components.
  • 3.
    For complex shapes involving multiple ”Sketch-Extrusion” steps, the model struggles to generate accurate results. It has difficulty remembering the previously drawn parts through the context, leading to errors.
  • 4.
    Due to the current work focusing only on basic instructions without a combined instruction set like drawing rectangles in one command, the CAD operation codes for 3D shapes tend to be lengthy. Moreover, due to contextual limitations of the model, especially with OpenECAD 0.55B and 0.89B using OpenELM which supports only 2048 tokens, and considering that image inputs also consume tokens, there are situations where the code cannot be completed.
6.5 Future Work
6.5.1 Refinement
Based on the above analysis, our future work will focus on the following areas:
  • 1.
    Enhancing the dataset involves several key strategies: expanding support for additional CAD instructions, annotating critical shape dimensions in input images to aid accurate size-based shape drawing by the model, selecting optimal angles for input images to showcase all features comprehensively, and employing diverse rendering tools to enhance the model’s ability to recognize shapes under varied lighting conditions and artistic styles.
  • 2.
    Improving the visual model is crucial as the current resolution limits its ability to distinguish between certain shapes. Therefore, training a visual model capable of handling higher resolutions is necessary. Additionally, integrating an image segmentation module like FastSAM[38] before the visual model can aid in identifying which parts compose the 3D shapes.
  • 3.
    Improving the language model’s capabilities is essential, as the current model supports a limited context length and has weak code generation abilities, resulting in incomplete or erroneous code. We can address this by using larger language models that support longer contexts, such as Phi-3-medium or LLaMA-3. Additionally, implementing Retrieval-Augmented Generation (RAG) [39] and prompting techniques can further enhance the model’s contextual understanding and code generation capabilities.
  • 4.
    Enhancing feedback integration with CAD tools. By adopting a conversational approach, partly generated CAD instructions can be promptly sent to the CAD tool for rendering. The rendering results can then be immediately returned to the model, helping it understand the current drawing status, identify potential errors, correct them, and continue generating code based on the target shape and the already rendered results.
At the same time, we have added a Constraints section to the dataset. We consider nine types of widely used constraint commands (see Table 5): making a line horizontal or vertical, fixing the size of a curve, making two points of curves coincident, making two lines parallel, perpendicular, or at a fixed angle, and making two curves tangent or mirrored. These constraints reduce the degrees of freedom in the sketch. A sketch with exactly zero degrees of freedom is fully defined, while one with fewer than zero degrees of freedom is over-constrained. We generated a 2D CAD dataset derived from the SketchGraphs dataset[18], resulting in a total of 628,499 2D sketches. The statistical distribution of the number of curves and constraints operations in the 2D dataset is shown in the Figure 10.
Command Name Parameters